# Bangkok 10330, Thailand

### v1

@@ 1 Introduction
It is well known that propagation of massless fields in de Sitter (dS) space is plagued with infrared divergences at distances larger than than the dS radius $H^{-1}$, with $H$ the Hubble constant fixing the expansion rate. In particular, the propagator of a massless minimally coupled scalar, like the inflaton during inflation, grows logarithmically Ford:1977in ; Ratra:1984yq ; Antoniadis:1985pj . These divergences are physical and show up in inflationary observables, as we found recently in the computation of the three-point function of scalar primordial fluctuations (bispectrum) at the next-to-leading order (NLO) in the slow-roll parameters of single field inflation, giving rise to sizeable corrections in the non-gaussianity parametrised by an observable parameter $f_{\text{NL}}$. In summary, we found two types of logarithmic corrections to the bispectrum Antoniadis:2024abm :
@@
$\big{\langle}\zeta\big{(}\vec{k}_{1}\big{)}\zeta\big{(}\vec{k}_{2}\big{)}\big{\rangle}$ $=(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}\big{)}\mathcal{P}_{k_{1}},$ (1)
$\big{\langle}\zeta\big{(}\vec{k}_{1}\big{)}\zeta\big{(}\vec{k}_{2}\big{)}\zeta\big{(}\vec{k}_{3}\big{)}\big{\rangle}$ $=(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{B}(k_{1},k_{2},k_{3}),$ (2)
where $\zeta$ is the gauge invariant scalar perturbation. The squeezed limit of the bispectrum is given by the well-known consistency relation Maldacena:2002vr ; Creminelli:2004yq
@@
$\qquad\qquad\qquad\quad\sim-(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}_{k_{1}}\left(3+k_{2}\frac{d}{dk_{2}}\right)\mathcal{P}_{k_{2}}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (3)
@@
$\big{\langle}\zeta\big{(}\vec{k}_{1}\big{)}\zeta\big{(}\vec{k}_{2}\big{)}\zeta\big{(}\vec{k}_{3}\big{)}\big{\rangle}_{k_{1}\rightarrow 0}\sim-n_{s}(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}_{k_{1}}\mathcal{P}_{k_{1}}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right].$ (4)
Note that all quantities above are assumed to be time independent (existence of large time limit) and the behaviour in the limit $k_{1}\rightarrow 0$ of the 3-point function to be the same as the behaviour of $\mathcal{P}_{k_{1}}$, so that the limit exists upon multiplication with $k_{1}^{3-n_{s}}$ ($k_{1}^{3}$ at lowest order), or equivalently $\lim_{k_{1}\to 0}(\mathcal{B}_{k_{1},k_{2},k_{3}}/\mathcal{P}_{k_{1}})$ is finite. Obviously, the existence of logarithmic divergences described above invalidates these assumptions and asks the question on the validity of the consistency relation in the squeezed limit, or its eventual modification when these divergences are taken into account. In this note, we study this question and we find that the consistency relation remains valid in the vicinity of these limits, at least to NLO where explicit expressions are known. We also show that the NLO expression of the bispectrum Antoniadis:2024abm is conserved in time.
Actually the squeezed limit of the bispectrum and the consistency relation up to NLO were studied in Renaux-Petel:2010paw using a different method based on taking the limit in the action, rather than directly of the correlator. Our analysis serves also as an independent check of our previous general result for the bispectrum at NLO.
@@
Before proceeding, we summarise our conventions. We consider the background inflationary metric to be $ds^{2}=-dt^{2}+a(t)^{2}d\vec{x}^{2}$ with $\vec{x}=(x^{1},x^{2},x^{3})$. The Hubble parameter is defined by $H=\dot{a}/a$, where the dot is $d/dt$, and the slow-roll parameters are defined by $\varepsilon_{1}=-\dot{H}/H^{2}$ and $\varepsilon_{i+1}=\dot{\varepsilon}_{i}/H\varepsilon_{i}$ ($i\geq 1$) [^2] In Antoniadis:2024abm , we called $\varepsilon_{i}$ the Hubble flow functions in order to emphasize the difference from the slow-roll parameters defined through an inflaton potential that will not appear in this paper.. Note that these parameters are functions of time. The conformal time $\tau$ is defined through $d\tau=dt/a(t)$.
@@ 2 The consistency relation
We start by briefly deriving the consistency relation Creminelli:2004yq ; Cheung:2007sv ; Creminelli:2011rh ; Creminelli:2012ed  [^3] See also Goldberger:2013rsa ; Pimentel:2013gza ; Kundu:2015xta ; Hui:2018cag for other derivations of the consistency relation and its generalisations.. In the case where one of the three momenta is much smaller than the others, say $k_{1}\ll k_{2},k_{3}$, it exits the horizon way before. Consequently, it is a frozen classical mode acting as the background for the other two. In order to compute the 3-point function in this limit, one can therefore multiply $\zeta\big{(}\vec{k}_{1}\big{)}$ by the 2-point function of $\zeta\big{(}\vec{k}_{2}\big{)}$ and $\zeta\big{(}\vec{k}_{3}\big{)}$ with $\zeta\big{(}\vec{k}_{1}\big{)}$ acting as a background and average over it. Namely we can write
@@
Actually, one can use a trick here. One can show that the long wavelength mode $\zeta_{L}\big{(}\vec{k}_{1}\big{)}$ can be seen as an adiabatic mode which can be added or subtracted by a change of coordinates Creminelli:2012ed ; Weinberg:2003sw . At leading order in $k_{1}$, it can be removed by a simple time-independent spatial dilatation. To see this, let us recall that the gauge has already been fixed here in the $\zeta$-gauge where the spatial distance is given by
$d\vec{\ell}\,^{2}=a^{2}(t)e^{2\zeta(t,\vec{x})}d\vec{x}^{2}.$ (6)
So performing a time-independent spatial dilatation [^4] Indeed the long wavelength mode can be considered as frozen. Its time dependence would be subleading in $k_{L}$. Moreover, we could also consider a special conformal transformation but its effect is also subleading. See Creminelli:2012ed for more details.
$\vec{x}\rightarrow e^{\zeta_{L}}\vec{x},$ (7)
@@
$d\vec{\ell}\,^{2}\rightarrow a^{2}e^{2\zeta(t,\vec{x})+2\zeta_{L}}d\vec{x}^{2}.$ (8)
@@
$=\big{\langle}\zeta(\vec{x}_{2})\zeta(\vec{x}_{3}\big{)}\big{\rangle}+\zeta_{L}(x_{B})\Big{(}\vec{x}_{2}\cdot\vec{\partial}_{\vec{x}_{2}}+\vec{x}_{3}\cdot\vec{\partial}_{\vec{x}_{3}}\Big{)}\big{\langle}\zeta(\vec{x}_{2})\zeta(\vec{x}_{3}\big{)}\big{\rangle}+\cdots,$ (9)
@@
$\vec{x}\rightarrow\vec{\tilde{x}}=\big{(}1+\zeta_{L}\big{)}\vec{x}.$ (10)
@@
The first term of (9) will give a vanishing contribution when one re-introduces it in (5) and averages over the long wavelength mode. Moreover, its position $\vec{x}_{B}$ was chosen so that $\zeta(\vec{x}_{B})$ provides a representative value of the long wavelength (namely almost constant) $\zeta_{L}$. Its choice is arbitrary and we are now interested in the second term of (9):
@@
$\begin{split}&\quad\zeta_{L}(x_{B})\Big{(}\vec{x}_{2}\cdot\vec{\partial}_{\vec{x}_{2}}+\vec{x}_{3}\cdot\vec{\partial}_{\vec{x}_{3}}\Big{)}\big{\langle}\zeta(x_{2})\zeta(x_{3}\big{)}\big{\rangle}\\ &=\int\frac{d^{3}\vec{k}_{B}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\zeta_{L}\big{(}\vec{k}_{B}\big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}\\ &\quad\times\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}e^{i\vec{x}_{B}\cdot\vec{k}_{B}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}\\ &=-\int\frac{d^{3}\vec{k}_{B}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}(2\pi)^{3}\zeta_{L}\big{(}\vec{k}_{B}\big{)}e^{i\vec{x}_{B}\cdot\vec{k}_{B}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}\\ &\quad\times\Big{[}6\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}+\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}\\ &\qquad\quad+\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\cdot\vec{\partial}_{\vec{k}_{2}+\vec{k}_{3}}\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\Big{]}\\ &=-\int\frac{d^{3}\vec{k}_{B}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}(2\pi)^{3}\zeta_{L}\big{(}\vec{k}_{B}\big{)}e^{i\vec{x}_{B}\cdot\vec{k}_{B}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}\\ &\quad\times\Big{[}3\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}+\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}+\cdots\Big{]},\end{split}$ (11)
@@
where the integration over $\vec{k}_{B}$ is around the origin $\vec{k}_{B}=0$ where $\zeta_{L}$ peaks in momentum space. In (11) we performed two integrations by part, and the dots in the last line denote vanishing terms as they are proportional to $(\vec{k}_{2}+\vec{k}_{3})\delta^{(3)}(\vec{k}_{2}+\vec{k}_{3})$. Moreover, it is important to notice that $\mathcal{P}(\vec{k}_{2},\vec{k}_{3})$ is the power spectrum without the momentum conservation $\vec{k}_{3}=-\vec{k}_{2}$ from $\delta^{(3)}(\vec{k}_{2}+\vec{k}_{3})$ applied. It is related to $\mathcal{P}_{k_{2}}$ as
$\mathcal{P}_{k_{2}}=\mathcal{P}\big{(}\vec{k}_{2},-\vec{k}_{2}\big{)}$ (12)
@@
$\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}=k_{2}\frac{d}{dk_{2}}\mathcal{P}_{k_{2}}$ (13)
@@
The last step is now to multiply (9) by $\zeta_{L}\big{(}\vec{k}_{1}\big{)}$ and to average over it. We are then left over only with the second term on the right hand side which was evaluated in (11). In Fourier space and after integrating over the momentum $\vec{k}_{B}$ we then obtain
$\quad\Big{\langle}\zeta_{L}\big{(}\vec{x}_{1}\big{)}\big{\langle}\zeta\big{(}\vec{x}_{2}\big{)}\zeta\big{(}\vec{x}_{3}\big{)}|\zeta_{L}\big{\rangle}\Big{\rangle}$
$=-\int\frac{d^{3}\vec{k}_{1}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}e^{i\vec{x}_{1}\cdot\vec{k}_{1}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}$
×$\qquad\qquad\qquad\qquad\qquad\times\mathcal{P}_{k_{1}}\left(3\mathcal{P}_{k_{2}}+k_{2}\frac{d\mathcal{P}_{k_{2}}}{dk_{2}}\right)\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (14)
@@
Equating this with (2) as claimed by (5), we obtain
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\rightarrow 0}=-\mathcal{P}_{k_{1}}\left(3\mathcal{P}_{k_{2}}+k_{2}\frac{d\mathcal{P}_{k_{2}}}{dk_{2}}\right)\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (15)

### v2

# Bispectrum at NLO in single field inflation: conservation and squeezed limit

@@ 1 Introduction
It is well known that propagation of massless fields in de Sitter (dS) space is plagued with infrared divergences at distances larger than than the dS radius $H^{-1}$, with $H$ the Hubble constant fixing the expansion rate. In particular, the propagator of a massless minimally coupled scalar, like the inflaton during inflation, grows logarithmically [1, 2, 3]. These divergences are physical and show up in inflationary observables, as we found recently in the computation of the three-point function of scalar primordial fluctuations (bispectrum) at the next-to-leading order (NLO) in the slow-roll parameters of single field inflation, giving rise to sizeable corrections in the non-gaussianity parametrised by an observable parameter $f_{\text{NL}}$. In summary, we found two types of logarithmic corrections to the bispectrum [4]:
@@
$\big{\langle}\zeta\big{(}\vec{k}_{1}\big{)}\zeta\big{(}\vec{k}_{2}\big{)}\big{\rangle}$ $=(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}\big{)}\mathcal{P}_{k_{1}},$ (1.1)
$\big{\langle}\zeta\big{(}\vec{k}_{1}\big{)}\zeta\big{(}\vec{k}_{2}\big{)}\zeta\big{(}\vec{k}_{3}\big{)}\big{\rangle}$ $=(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{B}(k_{1},k_{2},k_{3}),$ (1.2)
where $\zeta$ is the gauge invariant scalar perturbation. The squeezed limit of the bispectrum is given by the well-known consistency relation [5, 6]
@@
$\qquad\qquad\qquad\quad\sim-(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}_{k_{1}}\left(3+k_{2}\frac{d}{dk_{2}}\right)\mathcal{P}_{k_{2}}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (1.3)
@@
$\big{\langle}\zeta\big{(}\vec{k}_{1}\big{)}\zeta\big{(}\vec{k}_{2}\big{)}\zeta\big{(}\vec{k}_{3}\big{)}\big{\rangle}_{k_{1}\rightarrow 0}\sim-n_{s}(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}_{k_{1}}\mathcal{P}_{k_{1}}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right].$ (1.4)
Note that all quantities above are assumed to be time independent (existence of large time limit) and the behaviour in the limit $k_{1}\rightarrow 0$ of the 3-point function to be the same as the behaviour of $\mathcal{P}_{k_{1}}$, so that the limit exists upon multiplication with $k_{1}^{3-n_{s}}$ ($k_{1}^{3}$ at lowest order), or equivalently $\lim_{k_{1}\to 0}(\mathcal{B}_{k_{1},k_{2},k_{3}}/\mathcal{P}_{k_{1}})$ is finite. Obviously, the existence of logarithmic divergences described above invalidates these assumptions and asks the question on the validity of the consistency relation in the squeezed limit, or its eventual modification when these divergences are taken into account. In this note, we study this question and we find that the consistency relation remains valid in the vicinity of these limits, at least to NLO where explicit expressions are known. We also show that the NLO expression of the bispectrum [4] is conserved in time.
Actually the squeezed limit of the bispectrum and the consistency relation up to NLO were studied in [7] using a different method based on taking the limit in the action, rather than directly of the correlator. Our analysis serves also as an independent check of our previous general result for the bispectrum at NLO.
@@
Before proceeding, we summarise our conventions. We consider the background inflationary metric to be $ds^{2}=-dt^{2}+a(t)^{2}d\vec{x}^{2}$ with $\vec{x}=(x^{1},x^{2},x^{3})$. The Hubble parameter is defined by $H=\dot{a}/a$, where the dot is $d/dt$, and the slow-roll parameters are defined by $\varepsilon_{1}=-\dot{H}/H^{2}$ and $\varepsilon_{i+1}=\dot{\varepsilon}_{i}/H\varepsilon_{i}$ ($i\geq 1$) [^2] In [4], we called $\varepsilon_{i}$ the Hubble flow functions in order to emphasise the difference from the slow-roll parameters defined through an inflaton potential that will not appear in this paper.. Note that these parameters are functions of time. The conformal time $\tau$ is defined through $d\tau=dt/a(t)$.
@@ 2 The consistency relation
We start by briefly deriving the consistency relation [6, 8, 9, 10] [^3] See also [11, 12, 13, 14] for other derivations of the consistency relation and its generalisations.. In the case where one of the three momenta is much smaller than the others, say $k_{1}\ll k_{2},k_{3}$, it exits the horizon way before. Consequently, it is a frozen classical mode acting as the background for the other two. In order to compute the 3-point function in this limit, one can therefore multiply $\zeta\big{(}\vec{k}_{1}\big{)}$ by the 2-point function of $\zeta\big{(}\vec{k}_{2}\big{)}$ and $\zeta\big{(}\vec{k}_{3}\big{)}$ with $\zeta\big{(}\vec{k}_{1}\big{)}$ acting as a background and average over it. Namely we can write
@@
Actually, one can use a trick here. One can show that the long wavelength mode $\zeta_{L}\big{(}\vec{k}_{1}\big{)}$ can be seen as an adiabatic mode which can be added or subtracted by a change of coordinates [10, 15]. At leading order in $k_{1}$, it can be removed by a simple time-independent spatial dilatation. To see this, let us recall that the gauge has already been fixed here in the $\zeta$-gauge where the spatial distance is given by
$d\vec{\ell}\,^{2}=a^{2}(t)e^{2\zeta(t,\vec{x})}d\vec{x}^{2}.$ (2.2)
So performing a time-independent spatial dilatation [^4] Indeed the long wavelength mode can be considered as frozen. Its time dependence would be subleading in $k_{L}$. Moreover, we could also consider a special conformal transformation but its effect is also subleading. See [10] for more details.
$\vec{x}\rightarrow e^{\zeta_{L}}\vec{x},$ (2.3)
@@
$d\vec{\ell}\,^{2}\rightarrow a^{2}e^{2\zeta(t,\vec{x})+2\zeta_{L}}d\vec{x}^{2}.$ (2.4)
@@
$=\big{\langle}\zeta(\vec{x}_{2})\zeta(\vec{x}_{3}\big{)}\big{\rangle}+\zeta_{L}(x_{B})\Big{(}\vec{x}_{2}\cdot\vec{\partial}_{\vec{x}_{2}}+\vec{x}_{3}\cdot\vec{\partial}_{\vec{x}_{3}}\Big{)}\big{\langle}\zeta(\vec{x}_{2})\zeta(\vec{x}_{3}\big{)}\big{\rangle}+\cdots,$ (2.5)
@@
$\vec{x}\rightarrow\vec{\tilde{x}}=\big{(}1+\zeta_{L}\big{)}\vec{x}.$ (2.6)
@@
The first term of (2.5) will give a vanishing contribution when one re-introduces it in (2.1) and averages over the long wavelength mode. Moreover, its position $\vec{x}_{B}$ was chosen so that $\zeta(\vec{x}_{B})$ provides a representative value of the long wavelength (namely almost constant) $\zeta_{L}$. Its choice is arbitrary and we are now interested in the second term of (2.5):
@@
$\begin{split}&\quad\zeta_{L}(x_{B})\Big{(}\vec{x}_{2}\cdot\vec{\partial}_{\vec{x}_{2}}+\vec{x}_{3}\cdot\vec{\partial}_{\vec{x}_{3}}\Big{)}\big{\langle}\zeta(x_{2})\zeta(x_{3}\big{)}\big{\rangle}\\ &=\int\frac{d^{3}\vec{k}_{B}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\zeta_{L}\big{(}\vec{k}_{B}\big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}\\ &\qquad\times\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}e^{i\vec{x}_{B}\cdot\vec{k}_{B}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}\\ &=-\int\frac{d^{3}\vec{k}_{B}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}(2\pi)^{3}\zeta_{L}\big{(}\vec{k}_{B}\big{)}e^{i\vec{x}_{B}\cdot\vec{k}_{B}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}\\ &\qquad\quad\times\Big{[}6\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}+\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}\\ &\qquad\qquad\quad+\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\cdot\vec{\partial}_{\vec{k}_{2}+\vec{k}_{3}}\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\Big{]}\\ &=-\int\frac{d^{3}\vec{k}_{B}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}(2\pi)^{3}\zeta_{L}\big{(}\vec{k}_{B}\big{)}e^{i\vec{x}_{B}\cdot\vec{k}_{B}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}\\ &\quad\qquad\times\Big{[}3\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}+\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}+\cdots\Big{]},\end{split}$ (2.7)
@@
where the integration over $\vec{k}_{B}$ is around the origin $\vec{k}_{B}=0$ where $\zeta_{L}$ peaks in momentum space. In (2.7) we performed two integrations by part, and the dots in the last line denote vanishing terms as they are proportional to $(\vec{k}_{2}+\vec{k}_{3})\delta^{(3)}(\vec{k}_{2}+\vec{k}_{3})$. Moreover, it is important to notice that $\mathcal{P}(\vec{k}_{2},\vec{k}_{3})$ is the power spectrum without the momentum conservation $\vec{k}_{3}=-\vec{k}_{2}$ from $\delta^{(3)}(\vec{k}_{2}+\vec{k}_{3})$ applied. It is related to $\mathcal{P}_{k_{2}}$ as
$\mathcal{P}_{k_{2}}=\mathcal{P}\big{(}\vec{k}_{2},-\vec{k}_{2}\big{)}$ (2.8)
@@
$\Big{(}\vec{k}_{2}\cdot\vec{\partial}_{\vec{k}_{2}}+\vec{k}_{3}\cdot\vec{\partial}_{\vec{k}_{3}}\Big{)}\mathcal{P}\big{(}\vec{k}_{2},\vec{k}_{3}\big{)}=k_{2}\frac{d}{dk_{2}}\mathcal{P}_{k_{2}}$ (2.9)
@@
The last step is now to multiply (2.5) by $\zeta_{L}\big{(}\vec{k}_{1}\big{)}$ and to average over it. We are then left over only with the second term on the right hand side which was evaluated in (2.7). In Fourier space and after integrating over the momentum $\vec{k}_{B}$ we then obtain
$\Big{\langle}\zeta_{L}\big{(}\vec{x}_{1}\big{)}\big{\langle}\zeta\big{(}\vec{x}_{2}\big{)}\zeta\big{(}\vec{x}_{3}\big{)}|\zeta_{L}\big{\rangle}\Big{\rangle}$ $=-\int\frac{d^{3}\vec{k}_{1}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{2}}{(2\pi)^{3}}\frac{d^{3}\vec{k}_{3}}{(2\pi)^{3}}e^{i\vec{x}_{1}\cdot\vec{k}_{1}+i\vec{x}_{2}\cdot\vec{k}_{2}+i\vec{x}_{3}\cdot\vec{k}_{3}}(2\pi)^{3}\delta^{(3)}\big{(}\vec{k}_{2}+\vec{k}_{3}\big{)}$
×$\qquad\qquad\qquad\times\mathcal{P}_{k_{1}}\left(3\mathcal{P}_{k_{2}}+k_{2}\frac{d\mathcal{P}_{k_{2}}}{dk_{2}}\right)\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (2.10)
@@
Equating this with (2) as claimed by (2.1), we obtain
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\rightarrow 0}=-\mathcal{P}_{k_{1}}\left(3\mathcal{P}_{k_{2}}+k_{2}\frac{d\mathcal{P}_{k_{2}}}{dk_{2}}\right)\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (2.11)